p^2 - 5 is NEVER divisible by 3. WHY?

Grade:

1 Answers

ruchi yadav
10 years ago
Because P2 - 5 is sequence of ,

(1)2-5 = -4
(2)2-5 = -1
(3)2-5 = 4
(4)2-5 = 11
(5)2-5 = 20
(6)2-5 = 31
.........

So we can see all terms of this sequence are not divisible by 3.


Thank You
Ruchi
Askiitians Faculty